please help me solve this problem: the length of a rectangle is 4cm more than twice its width. the perimeter is 86cm. find the width and length.
Answer by stanbon(48568)

please help me solve this problem: the length of a rectangle is 4cm more than twice its width. the perimeter is 86cm. find the width and length.
----------------
Let the width be "x" cm.
Then the length is "x+4" cm
---------------
EQUATION:
2x + 2(x+4) = 86
x + x+4 = 43
2x = 39
x = 19.5 cm (this is the width)
x+4 = 23.5 cm ( this is the length)
==================
Cheers,
Stan H.